MOA-II Obsevation Log Monthly Summary 2006/11
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 4714-4749 | 36 | 177 | 10620 |
LMC | 709-781 | 73 | 449 | 134700 |
SMC | 250-260 | 11 | 71 | 21300 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 120 | 697 | 166620 (46h17m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
7 | 12 | 7 | 7 | 22 | 5 | 12 | 7 | 25 | 5 | 18 | 8 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
6 | 2 | 2 | 7 | 6 | 2 | 2 | 2 | 2 | 11 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
20 | 20 | 20 | 18 | 18 | 21 | 71 | 71 | 67 | 29 | 21 | 22 |
Field13 | Field14 | Field15 | Field16 |
25 | 26 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
11 | 10 | 8 | 7 | 7 | 7 | 7 | 7 | 7 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2006/11/1: thin cloud was illuminated by moon, sky was very bright, No good dark was taken after 11 tries. / night time = 26940(sec)
2006/11/2: use D1740-60, no good DARK for 300sec. / night time = 26700(sec)
2006/11/3: camera off for maintenance / night time = 26460(sec)
2006/11/4: Sockets for chip 2 and 7, O-rings, moolecular shieves are exchanged. Spare of telescope controller is tested. / night time = 26220(sec)
2006/11/5: Camera leak check and cool down. Image check. Spare controller test. Dome maintenance / night time = 25980(sec)
2006/11/6: Pressure test OK. Cooling, too slow. Chip1, socket bad contact / night time = 25740(sec)
2006/11/7: Camera warm-up. Swap the socket of chip1. Swap cryocooler. Signal check OK, Vacuum OK. Cooling OK. / night time = 25500(sec)
2006/11/8: Image check / night time = 25260(sec)
2006/11/9: maintainance finished. telescope analysis, RA, DEC, ROT offset adjustment, mirror washing / night time = 25020(sec)
2006/11/10: chi2 has noise. no good darks. / night time = 24780(sec)
2006/11/11: Flats(F1255 to F1264 are for gain, F1265 to F1301 are taken every 10 degree of rotator angle) / night time = 24540(sec)
2006/11/12: D1780(60sec) is OK. D1782(300sec) is slightly noisy on chip2 in "c". / night time = 24300(sec)
2006/11/13: / night time = 24000(sec)
2006/11/14: D1810(60sec), D1811(300sec) are OK. / night time = 23760(sec)
2006/11/15: D1813(60sec), D1814(300sec) are OK. / night time = 23580(sec)
2006/11/16: I adjusted (removed and inserted) the video board for chip2, because chip2 was noisy. After that chip2 had no strange noise, but bias level became low (about 150ADU). So I adjusted bias level (test118.lod, old is test115.lod). D1814(0sec) is OK. I take flats(F1303-1312) to estimate CCD gain. / night time = 23340(sec)
2006/11/17: D1845(15sec) is OK. F1314-1340 was taken every 10 deg of rotator angle. / night time = 23100(sec)
2006/11/18: D1847(300sec) is OK. / night time = 22860(sec)
2006/11/19: D1850(60sec), D1851(300sec) are OK. Galactic Bulge observation finished for this season. / night time = 22680(sec)
2006/11/20: Magellanic Bridge was observed by V & I filters. D1853(300sec), D1855(0sec) & D1856(10sec) are OK. / night time = 22380(sec)
2006/11/21: Pumping 17:30 to 21:00. D1861(300sec) is noisy on chip1, 2, but negligible. / night time = 22140(sec)
2006/11/22: Dark D1865(300sec), D1866(25sec) are OK. We took I-filter FLAT images. / night time = 21960(sec)
2006/11/23: No observation / night time = 21720(sec)
2006/11/24: Dark D1870(300sec) is OK. / night time = 21540(sec)
2006/11/25: No observation / night time = 21300(sec)
2006/11/26: Dark D1875(300sec) is OK. / night time = 21120(sec)
2006/11/27: I took the set of V-filter flat images(12sec). Dark D1881(12sec), D1882(300sec) are good. / night time = 20940(sec)
2006/11/28: No observation / night time = 20700(sec)
2006/11/29: No observation / night time = 20520(sec)
2006/11/30: No observation / night time = 20340(sec)
TOTAL NIGHT TIME: 705420(sec)
OPERATING RATE: 27.6 %
GB: 2.5 %
LMC: 21.6 %
SMC: 3.4 %
ALERT: 0.0 %
GRB: 0.0 %
Return to index
Previous month
Next month